OpenStudy (anonymous):

Find the center and radius of the sphere with equation x^2 + 3x + y^2 -4y + z^2 +12z = 36

terenzreignz (terenzreignz):

Okay... you need to master a little technique called... completing the square... you know of it? :D

OpenStudy (anonymous):

Yes, somewhat.

OpenStudy (radar):

\[x ^{2}+3x+___+y ^{2}-4y+___z ^{2}+12z+_____=36+_____+______+_____\] fill in completing the squares.

terenzreignz (terenzreignz):

Just in case you need a little review, here's an example of completing the square... \[\Large w^2+\color{red}4w =32\] the number in red (the one that is next to the variable WITH NO EXPONENT) that's the number you focus on. What you do is -GET HALF OF THAT NUMBER -AND THEN SQUARE THAT. To demonstrate, -half of 4 is 2 -square of 2 is 4 So, what we do now is add 4 to BOTH SIDES of the equation, like so \[\Large w^2 +4w \color{red}{+4}=32\color{red}{+4}\] You'll notice that the left side is not what you'd call a perfect square trinomial, illustrating, let's factor it, we get... \[\Large \color{blue}{(w+2)^2}=\color{red}{36}\] And that's completing the square, in a nutshell :D

terenzreignz (terenzreignz):

I hate critical typos, I meant "You'll notice that the left side is NOW what you'd call..."

OpenStudy (anonymous):

9/4, 4, 36 and then all of those values are also added to other side

terenzreignz (terenzreignz):

Impressive :) \[\large \left(x^2+3x + \frac94 \right)+(y^2 - 4y +4)+(z^2+12z+36)=36+\frac94+4+36\]

terenzreignz (terenzreignz):

You'll notice now that the left side now consists of three perfect square trinomials, right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I dont know how to do the first. the fraction messes me up

OpenStudy (anonymous):

But the others should be (y+2)^2 + (z+6)^2

terenzreignz (terenzreignz):

Okay, just a hint, remember the first thing you did with the coefficient of w? Half of it, that would be the constant term in each binomial. So with the x's, you had a 3/2 before you squared it... so... \[\large \left(x+\frac32\right)^2+(y -2)^2+(z+6)^2=\frac{313}4\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

Okay, yes.

terenzreignz (terenzreignz):

From here, it's just easy pickings. The general equation of a sphere is \[\Large (x-\color{red}h)^2 + (y -\color{blue}k)^2+(z-\color{green}l)^2 = \color{violet}r^2\] If you have the equation in this form, then your center is simply \[\Large (\color{red}h \ , \color{blue}k \ , \color{green}l)\] And your radius is simply \(\Large \color{violet}r\)

terenzreignz (terenzreignz):

So... can you now single out the values of h, k, l and r?

OpenStudy (anonymous):

(-3/2,2,-6)

OpenStudy (anonymous):

and radius should be Squareroot (313/4)

terenzreignz (terenzreignz):

or sqrt(313) / 2 :)

terenzreignz (terenzreignz):

Very nice though :) I can see you just needed a nudge in the right direction, didn't you? ;)

OpenStudy (anonymous):

Yes I did! Thank you for your help!!!

terenzreignz (terenzreignz):

No problem :)

OpenStudy (radar):

@terenzreignz good nudging.